OpenStudy (anonymous):

I got two different answers for the following: (w-9)/(w) - (9w-26)/(11w) Simplify. Would the correct answer be 1/11(2-(73/w) or 2/11 - 73/11w? Or are they both wrong?

OpenStudy (lalaly):

\[\frac{w-9}{w} \times \frac{11}{11} - \frac{9w-26}{11w}\]\[=\frac{11(w-9)-9w+26}{11w}\]

OpenStudy (lalaly):

\[\frac{11w-99-9w+26}{11w}\]\[=\frac{2w-73}{11w}\]

OpenStudy (anonymous):

oh well geeze, I was way off base then wasn't I?? LOL

OpenStudy (lalaly):

Lol yeh, if u hav another one try it on ur own and ill check it for u:D

OpenStudy (anonymous):

thank you, I am sure I will have.

OpenStudy (lalaly):

:D

OpenStudy (anonymous):

Here is one: for (2)/(sx^2) + (9)/(s^2x). The answer I got is (2s+9x)/(s^2x^2)

jhonyy9 (jhonyy9):

is right sure what you have got